OpenStudy (anonymous):

solve: sin^2a-cos^2a/sin^2a-sinacosa

OpenStudy (psymon):

Is all of sin^2a - sinacosa in the denominator, or only the sin^2a part?

OpenStudy (anonymous):

all of it

OpenStudy (campbell_st):

the numerator is the difference of 1 squares the denominator .. find the common factor.. then you'll see what will cancel from the numerator and denominator

OpenStudy (campbell_st):

oops difference of 2 squares...

OpenStudy (campbell_st):

its a question about factoring

OpenStudy (psymon):

\[\frac{ \sin^{2}a - \cos^{2}a }{ \sin^{2}a - sinacosa }\] My first instinct is to factor the top. When you have something minus something, no matter what those somethings are, you can factor it like a difference of squares: \[(x-a) = (\sqrt{x} - \sqrt{a})(\sqrt{x}+\sqrt{a})\] So thats what im going to do with the top.

OpenStudy (psymon):

\[\frac{ (sina-cosa)(sina+cosa) }{ \sin^{2}a-sinacosa }\]Now I can factor a sin out of the bottom: \[\frac{ (sina-cosa)(sina+cosa) }{ sina(sina-cosa) }\]Now you have a top and bottom factor that cancel: \[\frac{ sina + cosa }{ sina } = \frac{ sina }{ sina }+\frac{ cosa }{ sina }= 1 + cota\]

OpenStudy (anonymous):

aghh thank you. that actually helped a lot.

OpenStudy (psymon):

Yeah, sure :3